OpenStudy (geometry_hater):

Solve for x to the nearest degree. Solve for x to the nearest degree. @Mathematics

OpenStudy (geometry_hater):

i dont know what im doin

OpenStudy (anonymous):

use sine

OpenStudy (anonymous):

sin(x) = 15/18

OpenStudy (geometry_hater):

i get .83 this dosent make sense

OpenStudy (anonymous):

sin(x) = 5/6 x = ? You should be able to get that from the sine table.

OpenStudy (geometry_hater):

is it hypoth / adjacent ?

OpenStudy (anonymous):

Opposite side dividided by hypotenuse = sin(x).

OpenStudy (geometry_hater):

and its .83

OpenStudy (anonymous):

Correct. sin(x) is 0.83. Now, you need to find the value of x (according to the problem statement).

OpenStudy (anonymous):

5/6 = 0.833333

OpenStudy (geometry_hater):

its either 34 50 56 or 40 degrees

OpenStudy (anonymous):

It is 56 (based on calculator). :-)

OpenStudy (geometry_hater):

did you multiply something?

OpenStudy (anonymous):

No. I just opened the calculator and asked it what is the value of sin(56). It said 0.829 something.

OpenStudy (geometry_hater):

oh ok i see what you did now

OpenStudy (geometry_hater):

thank you

OpenStudy (anonymous):

You should use a table of values for the sine curve to look up, or a calculator. No easy way to memorize all possible values of sine.